Olmo, Inc., manufactures and sells two products: Product K0 and Product H9. The annual production and sales of Product of K0 is 700 units and of Product H9 is 700 units. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Activity Cost Pools Activity Measures Estimated Overhead Cost Expected Activity
Product K0 Product H9 Total
Labor-related DLHs $ 549,308 5,600 2,800 8,400
Production orders orders 52,319 800 400 1,200
Order size MHs 834,916 3,000 3,300 6,300
$ 1,436,543

The overhead applied to each unit of Product H9 under activity-based costing is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

A. $261.56 per unit

B. $911.26 per unit

C. $624.78 per unit

D. $513.05 per unit
Answer #1

Calculate overhead applied of each unit

Product H9
Labor-related 549308/8400*2800 = 183092
Production order 52319/1200*400 = 17440
Order size 834916/6300*3300 = 437349
Total 637881
Unit 700
Overhead cost per unit 911.26

So answer is b) $911.26